The problem is: One number is 8 more than the second number. Twice the second number is one more than the first. Find the numbers.
Answer by oberobic(2304)

x = first number
y = second number
x = 8y :: the first number = 8 times the second number
2y = x+1 :: twice the second number is one more than the first number
.
substitute x = 8y...
.
2y = 8y + 1
-6y = 1
y = 1/(-6) = -1/6
.
x = 8y
x = 8(-1/6) = -8/6 = -4/3
.
Checking...is twice the second number = 1 more than first?
.
2*y = 2(-1/6) = -2/6 = -1/3
x+1 = -4/3 + 1 = -1/3
Correct
.
Answer:
The numbers are x and y:
x = -4/3, and
y = -1/6.
.
Done.